# FactLens: Benchmarking Fine-Grained Fact Verification

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive capability in language generation and understanding, but their tendency to hallucinate and produce factually incorrect information remains a key limitation. To verify LLM-generated contents and claims from other sources, traditional verification approaches often rely on holistic models that assign a single factuality label to complex claims, potentially obscuring nuanced errors. In this paper, we advocate for a shift toward fine-grained verification, where complex claims are broken down into smaller sub-claims for individual verification, allowing for more precise identification of inaccuracies, improved transparency, and reduced ambiguity in evidence retrieval. However, generating sub-claims poses challenges, such as maintaining context and ensuring semantic equivalence with respect to the original claim. We introduce FactLens111link to our data and code redacted for review , a benchmark for evaluating fine-grained fact verification, with metrics and automated evaluators of sub-claim quality. The benchmark data is manually curated to ensure high-quality ground truth. Our results show alignment between automated FactLens evaluators and human judgments, and we discuss the impact of sub-claim characteristics on the overall verification performance.
@@
Kushan Mitra, Dan Zhang, Sajjadur Rahman, Estevam Hruschka Megagon Labs {kushan, dan_z, sajjadur, estevam}@megagon.ai
@@ 1 Introduction
Previous works on building fact-checking benchmarks focus on generating claims with a ground truth label, and in some cases provide the evidence/context to verify the claim, Aly et al. (2021); Schlichtkrull et al. (2023). Claims are generated using human annotators Aly et al. (2021), synthetic processes Bayat et al. (2023); Tang et al. (2024), or considering LLM outputs on Question-Answering tasks Wang et al. (2023). To increase the complexity of the fact-checking process, the claims are generated from source data of multiple domains & modalities, such as Wikipedia text and/or tables Thorne et al. (2018); Chen et al. (2019); Aly et al. (2021), Web Pages Schlichtkrull et al. (2023), Knowledge Graphs Kim et al. (2023), online posts/chats Wang et al. (2023); Li et al. (2023), and Q-A tasks from various domains such as statistics, finance, legal, etc Jacovi et al. (2024a).
These works also provide baseline fact-checking pipelines, which typically involves two main stages: (1) the retrieval of relevant evidence using Search APIs and multimodal data-lakes Tang et al. (2023); Schlichtkrull et al. (2023) and (2) the verification of claims based on that evidence using NLI-based, LLM-based and fine-tuned fact-verification models Li et al. (2023). Some works also explore delegating these steps entirely to an LLM-based policy framework Li et al. (2023); Peng et al. (2023).
Despite this structured pipeline, most existing methods rely on a holistic verification model, where complex claims are assigned a single factuality label, often obscuring the nuanced nature of the errors or inaccuracies in the claims. In this work, we echo the sentiments of Wang et al. (2023); Liu et al. (2019); Pan et al. (2023); Min et al. (2023); Si et al. (2024) for a shift towards fine-grained verification of complex claims, where claims are decomposed into smaller, more manageable sub-claims that can be individually verified. We additionally emphasise on the need to provide evaluation metrics to benchmark such fine-grained verification and enrich existing benchmarks with fine-grained verification labels.
@@
Achieving fine-grained verification, however, presents its own challenges. Decomposing a raw, complex claim into smaller sub-claims is not simply a matter of splitting it into sentences. Poorly constructed sub-claims can introduce a variety of issues: they may lose the context necessary for proper verification, lack atomicity, or misrepresent the original information by either omitting key details or introducing new ones. Ensuring the quality and verifiability of these sub-claims is therefore critical for the overall success of the verification process.
@@ 2.1 Evaluation Metrics
To perform fine-grained verification, each sub-claim needs to be independently verifiable. This requires the sub-claims to be properly contextualized to avoid any added ambiguity. Sufficiency measures whether the sub-claim is unambiguous and sufficiently contextualized with respect to the original claim. For example, in the original claim “Amanda Bauer attended the University of Cincinnati. The school’s nickname is Bearcats.” , the sub-claim “The school’s nickname is Bearcats” would be considered low in sufficiency because the reference to school was omitted in the decomposition, making it ambiguous.
@@
Ideal metric values: For an ideal claim decomposition, we expect the sub-claim to possess high atomicity, high sufficiency, high coverage and high readability while having low fabrication and low redundancy.
@@ 2.2 FactLens Evaluator
3 FactLens: Benchmarking Fine-grained Verification
3.1 Dataset Creation
The FactLens benchmark contains a dataset with ground-truth sub-claims and fine-grained labels. We use 733 instances from CoverBench Jacovi et al. (2024b), a fact-checking benchmark focused on complex claim verification sampled from diverse sources and domains, as the original claims.
The next step in fine-grained verification involves using a ‘verifier’ model to fact-check each sub-claim against the provided evidence. In this work, we use GPT-4o mini as our verifier model across all experiments to maintain consistency.
This verification method enables an enriched fine-grained evaluation. To compare the performance of fine-grained verification labels (for each sub-claim) with the holistic verification label, we aggregate the fine-grained labels as false if at least one of the fine-grained labels is also false; otherwise, the label is considered true.
3.2 Claim Decomposition: Model Performance
@@
We utilize the prompt defined in Table 3 to decompose claims using GPT-4o and Llama-3.1(405B). In Table 2 we tabulate the evaluation performance of both these models on the claim decomposition task using FactLens Evaluator.
@@ 4 Conclusion
5 Limitations
In this section, we note some limitations of our work.
5.1 Computation of Metrics
@@
To mitigate inconsistency, we provided specific instructions in the prompt (Table 4) to LLMs. we measured the agreement & correlation scores of LLMs with human judgement scores, observing fair-moderate agreement across most metrics. Furthermore, we propose our own definitions for computation of the FactLens Evaluator scores.
However, we acknowledge the limitations in our computational approach as well, with it relying on the method for entity extraction, which may produce variable results. We aim to propose more concrete definitions for these metrics in future works.
5.2 Evidence Retrieval
@@
5.3 Fact Verification Models
Previous works compared different verification models in the fact checking task. Tang et al. (2024) contrast the performance of MNLI-based, LLM-based and their proposed fine-tuned models for fact-verification. In this work, we choose to use GPT-4o mini as our only verifier model, as our aim is not to propose stronger models for verification; but to illustrate the benefits of fine-grained verification even using simpler off-the-shelf verifier models.
5.4 What are facts?
@@
Moreover, factual claims in real world-scenarios can often be temporal and domain-dependent in nature. For example, a claim such as “The legal drinking age is 18” is false in a country such as the United States, however may be true in the United Kingdom, indicating domain dependence. Similarly, evidence retrieved for the claim “Barack Obama is the President of the US” is temporal in nature.
We note that our results and experiments are also based on existing benchmarks, which do not account for all real-world scenarios.
@@ Appendix A Claim Decomposition
For the few-shot demonstrations we sample 4 instances from the FEVEROUS dataset ensuring no overlap with the CoverBench data. Finally in the prompt, we randomly select 3 of the 4 demonstrations and shuffle the order, to ensure there is no bias.
@@ B.1 Prompt for Evaluating Claim Decomposition
In Table 4 we provide the prompt to LLMs used for evaluating the claim decompositions across the 6 metrics defined in Section 2.1. We provide clear instructions using which LLMs can judge the claim decompositions across different dimensions. For all metrics except “coverage” and “redundancy”, the sub-claims are passed one at a time to obtain sub-claim level evaluation. “Coverage” and “redundancy” are used to judge the sub-claims as a whole, hence for these metrics, we provide the entire set of sub-claim decompositions for that instance. For “atomicity" we ask the LLM to output a label as per specific instructions, while for each of the remaining metrics we prompt the LLM to judge the instance with a “low", “medium” or “high” score.
We utilize the smaller and cheaper OpenAI model GPT-4o mini for this task, with temperature = 0.
@@ B.4 Agreement Scores on Synthetic Data
In Table 1 we observed fair to moderate correlation between humans and FactLens Evaluator scores through Pearson and Spearman correlation scores. We provide the corresponding p-values in Table 5, from which we conclude the correlation scores for atomicity, coverage, fabrication (LLM-Human) and redundancy (LLM-Human) are statistically significant.
@@
LLM 4e-10 5e-10 0.1402 0.4367 5.1e-4 2.63e-7 2.9e-4 2.4e-4 3.26e-6 7.78e-6
Statistical 1e-9 1e-9 0.1656 0.1764 7.7e-8 1.1e-6 0.4043 0.3316
Table 5: P-values from Table 1
@@ Appendix C Expert Annotations
To measure correlation between the FactLens Evaluator scores and human scores, we do an average of FactLens Evaluator score on a metric with both the annotators, and repeat for all metrics.
The human annotators were also used to generate ground-truth sub-claims as mentioned in Section 3.1.
@@ Appendix D Fine-Grained Verification
In order to perform verification, we utilize GPT-4o mini to judge if a claim is true or false based on the evidence provided. We obtain the ground truth evidence present in the CoverBench dataset.
@@ Appendix E Impact of Sub Claim Quality on Verification
From Figure 4, we observe that fabrication has the highest weight in magnitude, implying most influence in predicting whether the final label matches with ground truth. We see a negative weight for fabrication which is expected as lower fabrication indicates better quality sub-claims which in turn should have a positive effect on verification. Atomicity, sufficiency and coverage show a positive weight, as highly atomic, highly sufficient and high coverage sub-claims are expected to have a positive influence on verification.
@@ Appendix F FactLens Dataset Characteristics
Figure 5: Distribution of Number of Sub Claims in FactLens Benchmark

### v3

@@ Abstract
Large Language Models (LLMs) have shown impressive capability in language generation and understanding, but their tendency to hallucinate and produce factually incorrect information remains a key limitation. To verify LLM-generated contents and claims from other sources, traditional verification approaches often rely on holistic models that assign a single factuality label to complex claims, potentially obscuring nuanced errors. In this paper, we advocate for a shift towards fine-grained verification, where complex claims are broken down into smaller sub-claims for individual verification, allowing for more precise identification of inaccuracies, improved transparency, and reduced ambiguity in evidence retrieval. However, generating sub-claims poses challenges, such as maintaining context and ensuring semantic equivalence with respect to the original claim. We introduce FactLens [^1] https://github.com/megagonlabs/factlens , a benchmark for evaluating fine-grained fact verification, with metrics and automated evaluators of sub-claim quality. The benchmark data is manually curated to ensure high-quality ground truth. Our results show alignment between automated FactLens evaluators and human judgments, and we discuss the impact of sub-claim characteristics on the overall verification performance.
@@
Kushan Mitra Megagon Labs, USA kushan@megagon.ai                        Dan Zhang Megagon Labs, USA dan_z@megagon.ai
Sajjadur Rahmanthanks: Work done while at Megagon Labs. Adobe Inc. sajjadurr@adobe.com                        Estevam Hruschka Megagon Labs, USA estevam@megagon.ai
@@ 1 Introduction
Previous works on building fact-checking benchmarks focus on generating claims with a ground truth label, and in some cases provide the evidence/context to verify the claim Aly et al. (2021); Schlichtkrull et al. (2024). Claims are generated using human annotators Aly et al. (2021), synthetic processes Fatahi Bayat et al. (2023); Tang et al. (2024a), or considering LLM outputs on Question-Answering tasks Wang et al. (2024). To increase the complexity of the fact-checking process, the claims are generated from source data of multiple domains & modalities, such as Wikipedia text and/or tables Thorne et al. (2018); Chen et al. (2020); Aly et al. (2021), Web Pages Schlichtkrull et al. (2024), Knowledge Graphs Kim et al. (2023), online posts/chats Wang et al. (2024); Li et al. (2024), and QA tasks from various domains such as statistics, finance, legal, etc Jacovi et al. (2024).
These works also provide baseline fact-checking pipelines, which typically involves two main stages: (1) the retrieval of relevant evidence using Search APIs and multimodal data-lakes Tang et al. (2024b); Schlichtkrull et al. (2024) and (2) the verification of claims based on that evidence using NLI-based, LLM-based and fine-tuned fact-verification models Li et al. (2024). Some works also explore delegating these steps entirely to an LLM-based policy framework Li et al. (2024); Peng et al. (2023).
Despite this structured pipeline, most existing methods rely on a holistic verification model, where complex claims are assigned a single factuality label, often obscuring the nuanced nature of the errors or inaccuracies in the claims. In this work, we echo the sentiments of Wang et al. (2024); Liu et al. (2020); Pan et al. (2023); Min et al. (2023); Si et al. (2024) for a shift towards fine-grained verification of complex claims, where claims are decomposed into smaller, more manageable sub-claims that can be individually verified. We additionally emphasise on the need to provide evaluation metrics to benchmark such fine-grained verification, and enrich existing benchmarks with fine-grained verification labels.
@@
Achieving fine-grained verification, however, presents its own challenges. Decomposing a raw, complex claim into smaller sub-claims is not simply a matter of splitting it into sentences. Poorly constructed sub-claims can introduce a variety of issues: they may lose the context necessary for proper verification, lack atomicity, or misrepresent the original information by either omitting key details or introducing new fabricated ones. Ensuring the quality and verifiability of these sub-claims is, therefore, critical for the overall success of the verification process.
@@ 2.1 Evaluation Metrics
To perform fine-grained verification, each sub-claim needs to be independently verifiable. This requires the sub-claims to be properly contextualized to avoid any added ambiguity. Sufficiency measures whether the sub-claim is unambiguous and sufficiently contextualized with respect to the original claim. For example, in the original claim “Amanda Bauer attended the University of Cincinnati. The school’s nickname is Bearcats.” , the sub-claim “The school’s nickname is Bearcats” would be considered low in sufficiency because the reference to the school was omitted in the decomposition, making it ambiguous.
@@
Ideal metric values: For an ideal claim decomposition, we expect the sub-claim to possess high atomicity, high sufficiency, high coverage and high readability, while having low fabrication and low redundancy.
@@ 2.2 FactLens Evaluator
3 FactLens: Benchmarking Fine-grained Verification
3.1 Dataset Creation
The FactLens benchmark contains a dataset with ground-truth sub-claims and fine-grained labels. We use 733 instances from CoverBench Jacovi et al. (2024), a fact-checking benchmark focused on complex claim verification sampled from diverse sources and domains, as the original claims.
The next step in fine-grained verification involves using a ‘verifier’ model to fact-check each sub-claim against the provided evidence. In this work, we use GPT-4o-mini as our verifier model across all experiments to maintain consistency.
This verification method enables an enriched fine-grained evaluation. To compare the performance of fine-grained verification labels (for each sub-claim) with the holistic verification label, we aggregate the fine-grained labels as false if at least one of the fine-grained labels is also false; otherwise, the claim is considered true.
3.2 Claim Decomposition: Model Performance
We utilize the prompt defined in Table 3 to decompose claims using GPT-4o and Llama-3.1(405B). In Table 2, we tabulate the evaluation performance of both these models on the claim decomposition task using FactLens Evaluator.
@@ 4 Conclusion
Limitations
Computation of Metrics
@@
To mitigate inconsistency, we provided specific instructions in the prompt (Table 4) to LLMs. We measured the agreement & correlation scores of LLMs with human judgement scores, observing fair-moderate agreement across most metrics. Furthermore, we propose our own definitions for computation of the FactLens Evaluator scores. However, we acknowledge the limitations in our computational approach as well, with it relying on the method for entity extraction, which may produce variable results. We aim to propose more concrete definitions for these metrics in future works.
Evidence Retrieval
@@
Fact Verification Models
Previous works compared different verification models in the fact checking task. Tang et al. (2024a) contrast the performance of MNLI-based, LLM-based and their proposed fine-tuned models for fact-verification. In this work, we choose to use GPT-4o-mini as our only verifier model, as our aim is not to propose stronger models for verification; but to illustrate the benefits of fine-grained verification even using simpler off-the-shelf verifier models.
What are Facts?
@@
Moreover, factual claims in real world-scenarios can often be temporal and domain-dependent in nature. For example, a claim such as “The legal drinking age is 18” is false in a country such as the United States, however may be true in the United Kingdom, indicating domain dependence. Similarly, evidence retrieved for the claim “Barack Obama is the President of the US” is temporal in nature. We note that our results and experiments are also based on existing benchmarks, which do not account for all real-world scenarios.
@@ Appendix A Claim Decomposition
For the few-shot demonstrations we sample 4 instances from the FEVEROUS dataset Aly et al. (2021), ensuring no overlap with the CoverBench data. Finally in the prompt, we randomly select 3 of the 4 demonstrations and shuffle the order, to ensure there is no bias.
@@ B.1 Prompt for Evaluating Claim Decomposition
In Table 4, we provide the prompt to LLMs used for evaluating the claim decompositions across the 6 metrics defined in Section 2.1. We provide clear instructions using which LLMs can judge the claim decompositions across different dimensions. For all metrics except “coverage” and “redundancy”, the sub-claims are passed one at a time to obtain sub-claim level evaluation. “Coverage” and “redundancy” are used to judge the sub-claims as a whole, hence for these metrics, we provide the entire set of sub-claim decompositions for that instance. For “atomicity" we ask the LLM to output a label as per specific instructions, while for each of the remaining metrics we prompt the LLM to judge the instance with a “low", “medium” or “high” score.
We utilize the smaller and cheaper OpenAI model GPT-4o-mini for this task, with temperature = 0.
@@ B.4 Agreement Scores on Synthetic Data
In Table 1 we observed fair to moderate correlation between humans and FactLens Evaluator scores through Pearson and Spearman correlation scores. We provide the corresponding p-values in Table 5, from which we conclude the correlation scores for atomicity, coverage (both LLM-Human and Statistical-Human), fabrication, and redundancy (LLM-Human) are statistically significant.
@@
LLM 4e-10 5e-10 0.140 0.437 5.1e-4 2.63e-7 2.9e-4 2.4e-4 3.26e-6 7.78e-6
Statistical 1e-9 1e-9 0.166 0.176 7.7e-8 1.1e-6 0.404 0.332
Table 5: P-values from Table 1: Correlation of FactLens Evaluator scores with Human annotations on synthetic data
@@ Appendix C Expert Annotations
To measure correlation between the FactLens Evaluator scores and human scores, we do an average of FactLens Evaluator score on a metric with both the annotators, and repeat for all metrics. The human annotators were also used to generate ground-truth sub-claims (Section 3.1).
@@ Appendix D Fine-Grained Verification
In order to perform verification, we utilize GPT-4o-mini to judge if a claim is true or false based on the evidence provided. We obtain the ground truth evidence present in the CoverBench dataset.
@@ Appendix E Impact of Sub Claim Quality on Verification
Figure 5: Distribution of Number of Sub Claims in FactLens Benchmark
From Figure 4, we observe that fabrication has the highest weight in magnitude, implying most influence in predicting whether the final label matches with ground truth. We see a negative weight for fabrication which is expected as lower fabrication indicates better quality sub-claims which in turn should have a positive effect on verification. Atomicity, sufficiency and coverage show a positive weight, as highly atomic, highly sufficient and high coverage sub-claims are expected to influence verification positively.